The proportions of the market value of the firm's assets financed via debt, common stock, and preferred stock are called the firm's _____________________.

q45
Answer #1

Answer : Capital structure weights

suppose value of assets = 100 and  debt = 20, common stock = 50 and preferred stock = 30

then the capital structure weights are :  debt = 20/100 =0.20 = 20%, common stock = 50/100 = 0.50 = 50% and preferred stock = 30/100 =0.30 = 30%

Company uses these capital structure weights to find out WACC.

WACC = sum (weights x cost of source of fund) [Thumbs up please]